# Constraining Statistical Isotropy using 21cm Power Spectrum and Bispectrum

### v1

@@ 5.1 Power Spectra Results
Here we have performed integration over $\phi$ variable and also in the denominator we’ve approximated $P_{21}^{D}(k,\mu,z)\approx P_{21}^{0}(k,\mu,z)$, as the dipolar anisotropy is considered to be small compared to $1$. Similarly, for the quadrupolar asymmetry parameter $g_{0}$, the Fisher matrix element is
$F^{Q}_{P}(z)=\frac{V_{\mathrm{sur}}}{16\pi^{3}}\int\mathrm{d}k\mathrm{d}\mu\left[\frac{(P^{0}_{21}(k,\mu,z)-P_{\mathrm{N}}(k,\mu,z))\,k\,g(k)\,\mathcal{P}_{2}(\mu)}{P_{21}^{0}(k,\mu,z)}\right]^{2}$ (5.4)
We now discuss constraints on dipolar and quadrupolar anisotropy parameters respectively in Figures 2 and 3, using power spectrum. In these plots we show the cumulative errors as a function of $z$, considering a foreground cut $k_{\mathrm{fg}}=0.01h\mathrm{Mpc}^{-1}$ and a bit less optimistic value $k_{\mathrm{fg}}=0.02h\mathrm{Mpc}^{-1}$ (see (2.7)). The plots exhibit the following features.
@@
    The cumulative error $\sigma$ for both dipolar and quadrupolar anisotropies at $z=1$ is $\sim 10^{-3}$ for all three surveys considered in this paper.
@@
    As expected, power spectrum with a larger value of $k_{\mathrm{fg}}=0.01h\mathrm{Mpc}^{-1}$ has better constraining power as compared to that with $k_{\mathrm{fg}}=0.02h\mathrm{Mpc}^{-1}$. This is because larger foreground cuts render a larger number of $k$ modes unusable.
@@
    Regardless of the parameter ($A_{0}$ or $g_{0}$) constrained or the scaling ($F(k)$ or $g(k)$) used, PUMA gives the best constraints in the common redshift range $0.78\leq z\leq 2.55$.
@@
    For the dipolar case, Figure 2, we find that the error doesn’t change significantly with the scaling. Further, the difference between errors, for the chosen values of $k_{\mathrm{fg}}$ ($0.01$ or $0.02$) isn’t very significant. For a given redshift and scaling, the error increases as $n$ increases.
@@
    But we see a slightly different behaviour in case of quadrupolar anisotropy, Figure 3. We first notice that the difference between constraints obtained using with and without foregrounds keep decreasing as $n$ is increased on $-2$ to $2$. Apart from this, there isn’t any specific pattern in error obtained for a given survey and redshift by changing $n$.
Figure 5: Same as in Fig 3 but for bispectrum.
@@ 5.2 Bispectra Results
with $s_{123}=6,2,1$ for equilateral, isosceles, and scalene triangles respectively, $k_{\mathrm{f}}$ represents the fundamental frequency of the survey as was described in the previous section. Further, due to the smallness of the anisotropies, we approximate $P_{21}^{\mathrm{D}}$ as $P_{21}^{0}$, (4.2), i.e. the total power spectrum including thermal noise. Additionally, for degenerate configurations, i.e., $k_{1}=k_{2}+k_{3}$, we multiply the bispectrum variance by a factor of 2 [79]. The Fisher matrix element for quadrupolar anisotropy parameter $g_{0}$ will have the derivative of (4.2) with respect to $g_{0}$.
We must point out here that the bispectrum results are sensitive to the chosen values of $\Delta\mu_{1}$, $\Delta\phi$ [80] used for integration. Following [57], in this study, we take $\Delta\mu=0.04$ and $\Delta\phi=\pi/25$. The constraints using bispectrum are given in Figs. 4 and 5 for dipolar and quadrupolar anisotropies. We find that the constraints on the anisotropy parameters using the bispectrum are of the same order as that of the power spectrum. Also, similar to the power spectrum, the bispectrum also shows that PUMA has better constraining power.
@@
[…]
@@
$1$ 3 7 6 7 6 13
$52$ 6 7 7 4 9
Table 3: Same as in Table 2 but for quadrupolar anisotropy parameter $g_{0}$.
@@ 6 Conclusion and Outlook
    Out of all the surveys we have used in the study, PUMA works the best in the common redshift range.
@@
    The cumulative percent change in the case of bispectrum, when the foreground cuts are changed is found to be larger as compared to power spectrum. Thus, the constraining power of bispectrum is more sensitive to the foregrounds. This is evident by the percentage change in cumulative error $\rho$ given in Tables 2 and 3 for dipolar and quadrupolar parameters.
@@ Acknowledgments
Rahul Kothari acknowledges computing facilities availed through IIT Mandi seed Grant IITM/SG/DIS-ROS-SPA/111. Bhuwan Joshi acknowledges HTRA PhD scholarship while this work was done. We are also very thankful to Roy Maartens, Heyang Long, Sukhdeep Singh Gill and Suman Majumdar for discussions.

### v2

@@ 5.1 Power Spectra Results
Here we have performed integration over $\phi$ variable and also in the denominator we’ve approximated $P_{21}^{D}(k,\mu,z)\approx P_{21}^{0}(k,\mu,z)$, as the dipolar anisotropy is considered to be small compared to $1$. For the quadrupolar asymmetry, we’ve 5 parameters to constrain, thus the Fisher matrix element $F^{Q}_{P,ij}$ is written as
$F^{Q}_{P,ij}=\frac{V_{\mathrm{sur}}}{16\pi^{3}}\int k^{2}\mathrm{d}k\mathrm{d}\mu\mathrm{d}\phi\frac{\partial P_{21}^{Q}(\mathbf{k},z)}{\partial g_{i}}\frac{\partial P_{21}^{Q}(\mathbf{k},z)}{\partial g_{j}}\frac{1}{(P^{Q}_{21}(\mathbf{k},z))^{2}}$ (5.4)
Using (4.9) in this equation, we get
$F^{Q}_{P,ij}=\frac{V_{\mathrm{sur}}}{16\pi^{3}}\int\mathrm{d}k\mathrm{d}\mu\mathrm{d}\phi\ \mathfrak{F}_{i}\mathfrak{F}_{j}\left[\frac{(P^{0}_{21}(k,\mu,z)-P_{\mathrm{N}}(k,\mu,z))\,k\,g(k)}{P_{21}^{0}(k,\mu,z)}\right]^{2}$ (5.5)
The $\phi$ dependence in this integral comes only from the $\mathfrak{F}_{i}$, see (3.9). An integration over $\phi$ renders all non-diagonal terms zero. However, due to the identical functional dependence $\mu$ in the pair $\mathfrak{F}_{1}$ and $\mathfrak{F}_{2}$ & $\mathfrak{F}_{3}$ and $\mathfrak{F}_{4}$, the corresponding Fisher integrals are equal. As a result, the parameters $g_{1}$ and $g_{2}$ (and similarly $g_{3}$ and $g_{4}$) receive identical constraints. Thus for the quadrupolar study using power spectrum, we report constraints only on three parameters: $g_{1}$, $g_{3}$ and $g_{5}$.
We show the power spectrum constraints on dipolar anisotropy parameter in Figure 2 and those on quadrupolar anisotropy in Figures 3 and 4. In all these plots, cumulative errors as a function of $z$, considering foreground cuts $k_{\mathrm{fg}}=0.01h\mathrm{Mpc}^{-1}$ and $0.02h\mathrm{Mpc}^{-1}$ (see (2.7)) have been considered. As we discussed before, in case of quadrupolar anisotropy, we’ve only 3 parameters ($g_{1}$, $g_{3}$ and $g_{5}$) to constrain. To avoid clutter, we first check the behaviour of $\sigma_{g_{i}}$ in three separate plots for all three surveys with $g(k)=1$ in Figure 3. We find that constraints on $g_{1}$, $g_{3}$ and $g_{5}$ are of the same order $\lesssim\mathcal{O}(10^{-3})$, but the best constraints are found on the parameter $g_{5}\propto g_{20}$ (see (3.6) for $g_{2M}$ definition). We find the same trend in case of other scalings $g(k)$, again we’ve not included those plots to avoid clutter. Our best constrained parameter out of $g_{1}$, $g_{3}$ and $g_{5}$ is $g_{5}$, thus in Figure 4, we show only $\sigma_{g_{5}}$. The plots exhibit the following features.
@@
    The cumulative error $\sigma$ for both dipolar ($A_{0}$) and best constrained parameter ($g_{5}\propto g_{20}$) of quadrupolar anisotropies at $z=1$ is $\sim 10^{-3}$ for all three surveys considered.
@@
    As expected, power spectrum with a smaller value of $k_{\mathrm{fg}}=0.01h\mathrm{Mpc}^{-1}$ has a better constraining power as compared to that with $k_{\mathrm{fg}}=0.02h\mathrm{Mpc}^{-1}$. This is because larger foreground cuts render a larger number of $k$ modes unusable.
@@
    Regardless of the parameter ($A_{0}$ or $g_{5}$) constrained or the scaling ($F(k)$ or $g(k)$) used, PUMA gives the best constraints in the common redshift range $0.78\leq z\leq 2.55$.
@@
    For the dipolar case, Figure 2, we find that the error doesn’t change significantly with the scaling. Further, the difference between errors, for the chosen values of $k_{\mathrm{fg}}$ ($0.01h$ or $0.02h$) isn’t very significant. For a given redshift and scaling, the error slightly increases as $n$ increases.
@@
    But we see a slightly different behaviour in case of quadrupolar anisotropy parameter $g_{5}$ in Figures 3 and 4. We first notice that the difference between constraints obtained with and without foregrounds keep decreasing as $n$ is increased on 2-2 to $2$. Apart from this, there isn’t any specific pattern in error obtained for a given survey and redshift by changing $n$.
@@ 5.2 Bispectra Results
with $s_{123}=6,2,1$ for equilateral, isosceles, and scalene triangles respectively, $k_{\mathrm{f}}$ represents the fundamental frequency of the survey as was described in the previous section. Further, due to the smallness of the anisotropies, we approximate $P_{21}^{\mathrm{D}}$ as $P_{21}^{0}$, (4.2), i.e. the total power spectrum including thermal noise. Additionally, for degenerate configurations, i.e., $k_{1}=k_{2}+k_{3}$, we multiply the bispectrum variance by a factor of 2 [84].
For the quadrupolar asymmetry, the Fisher matrix will be written as
$F^{Q}_{B,ij}(z)=\frac{1}{4\pi}\sum_{k_{1}k_{2}k_{3}}\int\mathrm{d}\mu_{1}\int_{0}^{2\pi}\mathrm{d}\phi\frac{\partial B_{21}^{Q}}{\partial g_{i}}\frac{\partial B_{21}^{Q}}{\partial g_{j}}\frac{1}{\Delta B^{2}}$ (5.8)
with $B_{21}^{Q}$ given in (4.16) and $\Delta B^{2}$ in (5.7) with $P^{\mathrm{Q}}_{21}$ instead of $P^{\mathrm{D}}_{21}$. In addition to functions $\mathfrak{F}_{i}$ having $\phi$ dependencies, the kernels $Z_{i}$ are also be $\phi$ dependent. Thus unlike power spectrum case Subsection 5.1, there won’t be any identical constraint on the parameters $g_{1}$ and $g_{2}$ (and similarly $g_{3}$ and $g_{4}$) this time and we need to constrain all 5 parameters.
Figure 6: Same as Fig 3 but for bispectrum. Again, we find that $g_{5}$ is the best constrained parameter.
Figure 7: Same as in Fig 4 but for bispectrum.
We must point out here that the bispectrum results are sensitive to the chosen values of $\Delta\mu_{1}$, $\Delta\phi$ [85] used for integration. Following [57], in this study, we take $\Delta\mu=0.04$ and $\Delta\phi=\pi/25$. We give the constraints on dipolar anisotropy using bispectrum in Figure 5, while quadrupolar anisotropy results are given in Figures 6 and 7. As was done in the case of power spectrum, in Figure 6, we first check the $\sigma_{g_{i}}$ trend for $g(k)=1$. We find that just like power spectrum, bispectrum gives best constraints on $g_{5}$. Thus for the rest of the scalings shown in Figure 7, we show cumulative error only for $g_{5}$. We find that the constraints on the anisotropy parameters using the bispectrum are of the same order as that of the power spectrum. Also, similar to the power spectrum, the bispectrum also shows that PUMA has better constraining power. This reflects the conventional understanding that distributing the total collecting area across many small interferometric elements, such as PUMA’s $\sim 32,000$ dishes, enhances sensitivity by broadening the field of view.
@@
where $\sigma_{x}(z)$ is the cumulative error obtained for redshift $\leq z$ with $k_{\mathrm{fg}}=x$. This quantity, for both dipolar and quadrupolar anisotropies is tabulated in Tables 2 and 3 respectively. From Table 2, it is clear that percent cumulative change $\rho$ for bispectrum for a given $z$ and survey is greater than or equal to that of power spectrum. E.g., for PUMA $\rho$ values obtained for bispectrum is 20% which means that when the foregrounds are changed from $0.02h$ to $0.01h$, the change is 20% while for power spectrum the change is merely 6%. The same pattern is seen in Table 3.
@@
11 1 9 4 5 4 14
$2.5$ 1 7 3 8 4 10
Table 3: Same as in Table 2 but for quadrupolar anisotropy parameter $g_{5}$ with $g(k)=1$.
@@ 6 Conclusion and Outlook
    Further, out of all quadrupolar anisotropy parameters, $g_{i}$, the best constraints are obtained on $g_{5}$.
@@
    Out of all the surveys we have used in the study, PUMA works the best in the common redshift range for both power and bispectrum.
  4. 4.
    The cumulative percent change in the case of bispectrum, when the foreground cuts are changed is found to be larger as compared to power spectrum. Thus, the constraining power of bispectrum is more sensitive to the foregrounds. This is evident by the percentage change in cumulative error $\rho$ given in Tables 2 and 3 for dipolar and quadrupolar parameters. We believe that this is due to the fact that the bispectrum for its description needs information of three Fourier modes in contrast to the power spectrum that needs only two. Thus, a bispectrum configuration depending upon three Fourier modes is more likely to be discarded on account of the foreground conditions as compared to the power spectrum, which needs only two modes. To test our hypothesis, we calculate the change in the number of allowed configurations for both the power spectrum and the bispectrum for HIRAX at $z=1$ when $k_{\mathrm{fg}}$ is changed from $0.01h\mathrm{Mpc}^{-1}$ to $0.02h\mathrm{Mpc}^{-1}$. The results are shown in Table 4. It can be seen that the percent change in the number of allowed configurations for bispectrum is $\sim 6$ times as large as that of power spectrum. This continues to be case for other cases for various surveys and redshifts.
$0.01h$ $0.02h$ % change
Power Spectrum 240,518 224,550 6.6-6.6
Bispectrum 8,271,008 5,166,738 37.5-37.5
Table 4: Total number of allowed configurations in power spectrum and bispectrum as a function of foreground cuts $k_{\mathrm{fg}}$. The numbers are shown for HIRAX at $z=1$.
@@ Acknowledgments
Rahul Kothari acknowledges computing facilities availed through IIT Mandi seed Grant IITM/SG/DIS-ROS-SPA/111. Bhuwan Joshi acknowledges HTRA PhD scholarship while this work was done. We are also very thankful to Roy Maartens, Heyang Long, Sukhdeep Singh Gill, Manoj Rana and Suman Majumdar for discussions. We are also very thankful to the anonymous referee, whose comments helped us considerably improve the quality of the paper.
Appendix A Calculation of azimuthal and polar angles
In this appendix, our objective is to calculate the azimuthal and polar angles of the three vectors $\mathbf{k}_{i}$ that form the triangle, i.e., $\sum_{i}\mathbf{k}_{i}=0$. Although, the expressions for azimuthal angles, i.e., $\mu_{2}$ and $\mu_{3}$ is a known result [87], but to the best of our knowledge there are no such results in the literature for polar angles. So for completeness sake, here we give a complete derivation of calculations of all the involved angles. We assume a general LoS direction and later specialize to our case by considering it along $\hat{z}$. We want to express a vector $\mathbf{x}$ in terms of known vectors $\bm{\alpha}$ and $\bm{\beta}$ given (a) the angle between $\mathbf{x}$ and $\bm{\alpha}$ is $\theta$ and (b) the angle $\phi\in[0,2\pi]$ between $\mathbf{n}=\bm{\alpha}\times\mathbf{x}$ and the projection of $\bm{\beta}$ in the plane formed by the vectors $\mathbf{n}$ and $\mathbf{n}\times\bm{\alpha}$. Thus,
$\cos\phi=\frac{\alpha}{\sqrt{R}}\left[\bm{\beta}-\frac{(\bm{\beta}\cdot\bm{\rho})\bm{\rho}}{\rho^{2}}\right]\cdot\frac{\mathbf{n}}{n},\quad R=\alpha^{2}\beta^{2}-(\bm{\alpha}\cdot\bm{\beta})^{2}$ (A.1)
where $\bm{\rho}$ is normal to the plane formed by $\mathbf{n}=\bm{\alpha}\times\mathbf{x}$ and $\mathbf{n}\times\bm{\alpha}$. Although, there can be two normal vectors, in our case we take
$\bm{\rho}=(\mathbf{n}\times\bm{\alpha})\times\mathbf{n}$ (A.2)
In order to express $\mathbf{x}$ in terms of the given vectors and angles $\theta$ and $\phi$, the first step is to write the vector $\mathbf{x}$ in terms of linear combination of right handed orthogonal vector basis, viz., $\{\bm{\alpha},\bm{\beta}\times\bm{\alpha},\bm{\alpha}\times(\bm{\beta}\times\bm{\alpha})\}$
$\mathbf{x}=a\bm{\alpha}+b(\bm{\beta}\times\bm{\alpha})+c(\bm{\alpha}\times(\bm{\beta}\times\bm{\alpha}))$ (A.3)
As the vectors are mutually orthogonal, the basis is linearly independent. Our goal now is to calculate these coefficients. The coefficient $a$ is easily calculated by taking dot product of (A.3) with $\bm{\alpha}$ and we find
$a=\frac{x\cos\theta}{\alpha}$ (A.4)
To calculate coefficients $b$ and $c$, we first notice that
$n^{2}=|\bm{\alpha}\times\mathbf{x}|^{2}=x^{2}\alpha^{2}\sin^{2}\theta=\alpha^{2}\left[c^{2}\alpha^{2}|\bm{\alpha}\times\bm{\beta}|^{2}+b^{2}R\right]$ (A.5)
Another equation for these coefficients is obtained from (A.1) which after plugging in expressions of $\bm{\rho}$ & $\mathbf{n}$ gives
$n^{2}\cos^{2}\phi=b^{2}\alpha^{2}R$ (A.6)
Now we have two equations (A.5) and (A.6), solving these two gives
$c=\frac{x\sin\theta\sin\phi}{\alpha\ |\bm{\alpha}\times\bm{\beta}|},\quad b=\frac{x\cos\phi\sin\theta}{\sqrt{R}}$ (A.7)
where we have only considered the positive roots to get results consistent with the literature. Now we can check our results for the particular case when $\bm{\alpha}=\mathbf{k}_{1}$, $\mathbf{x}=\mathbf{k}_{2}$ as two sides of the triangle and LoS along Z axis, i.e., $\bm{\beta}=\hat{z}$. Plugging in these and after some simplification we get our coefficients as
$a=\frac{k_{2}\cos\theta_{12}}{k_{1}},\ b=\frac{k_{2}\cos\phi\sin\theta_{12}}{k_{1}\sqrt{1-\mu^{2}}},\ c=\frac{k_{2}\sin\theta_{12}\sin\phi}{k_{1}^{2}\sqrt{1-\mu^{2}}}$ (A.8)
where $\mu$ is the angle between $\mathbf{k}_{1}$ and $\hat{z}$. Taking dot product of (A.3) with $\hat{z}$ thus gives $\mu_{2}$ which we find to be
$\mu_{2}=\mu\cos\theta_{12}+\sqrt{1-\mu^{2}}\sin\theta_{12}\sin\phi$ (A.9)
Using the fact that vectors $\mathbf{k}_{i}$ form a triangle, we calculate $\mu_{3}$ using $\mathbf{k}_{3}\cdot\hat{z}$. Now we can proceed to calculating the polar angles, to that end, we notice that the polar angle $\Phi$ for a given vector $\mathbf{X}$ is
$\tan\Phi=\frac{\mathbf{X}\cdot\hat{y}}{\mathbf{X}\cdot\hat{x}}$ (A.10)
Conventionally $\mathbf{k}_{1}$ is taken in the XZ plane so that $\phi_{1}=0$. So we only need to calculate $\phi_{2}$ and $\phi_{3}$ which we find
$\tan\phi_{2}$ $=\frac{\cos\phi\sin\theta_{12}}{\sqrt{1-\mu^{2}}\cos\theta_{12}-\mu\sin\theta_{12}\sin\phi},$ (A.11)
$\tan\phi_{3}$ $=\frac{k_{2}\cos\phi\sin\theta_{12}}{\sqrt{1-\mu^{2}}(k_{1}+k_{2}\cos\theta_{12})-\mu k_{2}\sin\theta_{12}\sin\phi}$ (A.12)